Two balloons are charged with a different quantity of charge: balloon A with 5 nC and balloon B with 4 nC. They are held apart at a separation distance of 3.5 meters. Determine the magnitude of the electrical force of repulsion between them.


1
Expert's answer
2022-02-28T10:38:56-0500

Given:

"q_1=5*10^{-9}\:\rm C"

"q_2=4*10^{-9}\:\rm C"

"r=3.5\:\rm m"


The Coulomb's law says



"F=k\frac{|q_1||q_2|}{r^2}""F=9*10^9*\frac{5*10^{-9}*4*10^{-9}}{3.5^2}=1.47*10^{-8} \:\rm N"